OpenStudy (kawii2004):

Which equation can be used to find the answer? After two months, Rachel has saved enough money to buy an MP3 player. During the second month, she saved $87 and she pays $199 for the MP3 player. How much did Rachel save during the first month? A. m = 87 • 2 m = 174 $174 B. 199 + 87 = m m = 286 $286 C. m + 87 = 199 m = 112 $112 D. m – 87 = 199 m = 287 $287

OpenStudy (anonymous):

C. because you know one month and the price but you dont know the first month so you would be adding the months and your gonna subtract the second month from the total to find the first month

OpenStudy (kawii2004):

okay thank you so much

OpenStudy (kawii2004):

Which equation can be used to find the answer? Scott works a total of 14.5 hours after school from Monday to Friday. After the weekend, he has worked a total of 27.75 hours for the entire week. How many hours does Scott work on the weekend? A. w = 14.5 + 27.75 w = 42.25 42.25 hours B. w = 14.5 + 27.75 w = 41.25 41.25 hours C. w + 14.5 = 27.75 w = 13.25 13.25 hours D. w – 14.5 = 27.75 w = 13.75 13.75 hours

OpenStudy (kawii2004):

i have another

OpenStudy (anonymous):

C. the same explaination as last problem

OpenStudy (kawii2004):

okay

OpenStudy (kawii2004):

Which equation can be used to find the answer? An auto part warehouse shipped 96 cases of oil filters and still had 172 cases on its shelves. How many cases did it have on its shelves before the shipment? A. c + 96 = 172 c = 76 76 cases B. c ÷ 2 = 96 c = 192 192 cases C. c ÷ 2 = 172 c = 344 344 cases D. c – 96 = 172 c = 268 268 cases

OpenStudy (anonymous):

D.

OpenStudy (kawii2004):

okay

OpenStudy (kawii2004):

Which equation can be used to find the answer? Camilla needs $65 for a pair of hiking shoes. She already has $19 saved. How much more does she need to save? A. n + 19 = 65 n = 46 $46 B. n ÷ 2 = 19 n = 38 $38 C. n – 19 = 65 n = 84 $84 D. n ÷ 2 = 65 n = 130 $130

OpenStudy (anonymous):

A.

OpenStudy (kawii2004):

Which equation can be used to find the answer? Clarice spent $27 of her savings on clothes and has $49 of savings left. How much did she have in savings before buying the clothes? A. m + 27 = 49 m = 22 $22 B. m – 27 = 49 m = 76 $76 C. m ÷ 2 = 27 m = 54 $54 D. m ÷ 2 = 49 m = 98 $98

OpenStudy (kawii2004):

this the last one

OpenStudy (anonymous):

B.

OpenStudy (kawii2004):

okay thank you

OpenStudy (anonymous):

no problem

OpenStudy (kawii2004):

you are smart

OpenStudy (anonymous):

what type of math is that

OpenStudy (kawii2004):

i thing ummm idk

OpenStudy (anonymous):

i think its algebra 1

geerky42 (geerky42):

Pre-algebra is my guess.

OpenStudy (kawii2004):

yeah wait dont leave i have some more

OpenStudy (anonymous):

true thats a good one im in trigonometry as a softmore

OpenStudy (kawii2004):

What value of n makes the equation true? n – 29.4 = 13.9 A. 15.5 B. 16.5 C. 42.3 D. 43.3

OpenStudy (anonymous):

C.

OpenStudy (kawii2004):

The perimeter of this shape is 82 cm. Irregular quadrilateral What is the length of side d? cm

OpenStudy (anonymous):

do you have any lengths

OpenStudy (kawii2004):

wait hold up

OpenStudy (kawii2004):